The atmospheric pressure (measured in millibars and given the symbol P) decreases with height (h, in miles) according to:
P = 29.92xe^(-h/5)
a) Find the pressure at a height of 4 miles.
b) At what height (in miles) is the atmospheric pressure 10 millibars?

Answer by lwsshak3(11628)

The atmospheric pressure (measured in millibars and given the symbol P) decreases with height (h, in miles) according to:
P = 29.92xe^(-h/5)
..
a) Find the pressure at a height of 4 miles.
P=29.92*e^(-4/5)≈13.44 millibars
..
b) At what height (in miles) is the atmospheric pressure 10 millibars?
10=29.92*e^(-h/5)
10/29.92=e^(-h/5)
take log of both sides
ln(10/29.92)=-h/5 lne
lne=1
-h=ln(10/29.92)*5≈-5.48
h=5.48
Height at which the atmospheric pressure is 10 millibars=5.48 miles